Circular No. 7579

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001R IN NGC 5172 D. Weisz, M. Modjaz, and W. D. Li, University of California at Berke-

ley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 18.5) on unfiltered images taken with the 0.8m Katzman Automatic Imaging Telescope (KAIT) on Jan. 6.5, Feb. 3.5, and 4.5 UT. SN 2001R is located at α = 13^{h}29^{m}16^{s}.93, δ = +17^{o}02′41′′.2 (equinox 2000.0), which is 33′′.2 west and 25′′.0 south of the nucleus of NGC 5172. A KAIT image taken on 2000 June 20.2 showed nothing at this position (limiting mag ∼ 19.0). SN 1998cc also occurred in NGC 5172.

SUPERNOVAE 2001A AND 2001M A. V. Filippenko, R. Chornock, and M. Modjaz, University of Cali-

fornia at Berkeley, report that inspection of CCD spectra (range 330–1000 nm) of SN 2001A (cf. IAUC 7554), obtained on Feb. 1 UT with the Shane 3-m reflector at Lick Observatory, reveals that the object is a type-Ia supernova, roughly 5–6 weeks past maximum brightness. Spectra of SN 2001M (taken in the same way at Lick) confirm that it is a type-Ic supernova, 1–2 weeks past maximum brightness (cf. IAUC 7576).

COMET C/2001 C1 (LINEAR) Additional observations and the following very uncertain parabolic or-

bital elements appear on MPEC 2001-C07:

T = 2002 Jan. 27.671 TT  ω = 210.^{o}683
Ω = 33.970 }2000.0
q = 5.51440 AU  i = 68.112
2001 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Jan. 31 14^{h}41.^{m}70 −10^{o}08.^{′}5 6.028 6.108 90.^{o}0 9.^{o}3 17.8
Feb. 5 14 41.87 −10 31.0 5.929 6.092  94.9 9.3 17.7
10 14 41.76 −10 52.9 5.830 6.077  99.8 9.2 17.7
15 14 41.36 −11 14.3 5.733 6.062 104.9 9.1 17.6
20 14 40.66 −11 35.2 5.638 6.048 110.0 8.8 17.6
25 14 39.64 −11 55.4 5.545 6.033 115.1 8.5 17.5
Mar. 2 14 38.30 −12 15.2 5.456 6.019 120.4 8.2 17.5
7 14 36.65 −12 34.4 5.371 6.004 125.7 7.7 17.4
12 14 34.68 −12 53.1 5.291 5.990 131.0 7.2 17.4
2001 February 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green